OpenStudy (calculusxy):

There is a "40% off sale" at a store. Erica wants to buy a rowboat. The sale price is $351. What was the original price?

OpenStudy (calculusxy):

@phi @amistre @radar @Darrius

OpenStudy (phi):

call the original price "x" x - 40% of x is how much you pay x - 40% of x = 351 can you find x ?

OpenStudy (calculusxy):

Is it also like is/of? I set up a proportion of 341/x=4/10

OpenStudy (phi):

if you are careful. x is 100% of the cost. the sale takes 40% off : 100% - 40% = 60% the sales price is 60% of the full cost. Let's change 60% to 0.6 we can say \[ \frac{x}{1} = \frac{351}{0.6} \] that says full price is to 1 as sales price is to 0.6

OpenStudy (phi):

If you do it this way x - 40% of x = 351 change 40% to 0.4 x - 0.4x = 351 "factor out" x x(1-0.4) = 351 1- 0.4 = 0.6 , so you get 0.6x = 351 divide both sides by 0.6 x = 351/0.6

OpenStudy (calculusxy):

Just a short question. What do you mean by "factoring out"?